Journalist Report – March 10th

Sol 26 – It’s not over until it’s over… but now it’s over!

“’I’ve always wanted to see a Martian,’ said Michael. ‘Where are they, Dad? You promised.’
‘There they are,’ said Dad, and he shifted Michael on his shoulder and pointed straight down.
The Martians were there. Timothy began to shiver.
The Martians were there—in the canal—reflected in the water.”

– Chapter 26 of The Martian Chronicles by Ray Bradbury

Waking up to our last Martian sunrise was the strangest feeling, like ignoring the elephant in the room. We were due to return to Earth today, and staying focused in these circumstances was not easy! Nevertheless, Corentin led us for our final workout session of the mission, and we had breakfast, almost as if today was a normal day.

An EVA crew composed of Quentin, Alexandre, and Corentin then proceeded to perform the last EVA, in order to disassemble and bring back to the Hab all the atmospheric instruments deployed over the course of the mission: the LOAC, the field mill, Purple Air, the weather station, and the Mega Ares antenna. To be ready for any potential issue, Quentin had requested a 4-hour EVA for redundancy, but the crew ended up completing their objectives in under two hours. After a tear-jerking speech by Quentin over the radio and a few last photos with the station as a backdrop, the crew headed back to the station with their trunks full.
Performing my HabCom duties, I was half-listening to what Jérémy and Adrien were discussing while doing the dishes, and one question caught my attention – what if we had made different choices? In the sense: what if our paths had led us elsewhere than to where we are now? The very thought of there being a life in which I would not have lived through this experience threw me off completely, and for the first time, I felt that every choice I had made in my life had contributed to leading me here.

After lunch, we were “surprised” by a second emergency protocol simulation, during which we detected a breach in the GreenHab. Added difficulty: as planned by Quentin and Alexandre, a crewmember pretended to feel ill in the Science Dome during the sensor checks to disrupt the execution of the protocol. Despite it, the breach was sealed in time and the “ill” crewmember was safely returned to the Hab by our Health and Safety Officer. We continued finalizing and “disassembling” our experiments throughout the day; Adrien and Corentin took apart the aquaponic system, but not before they harvested all the greens it had produced! A feast awaits tonight, after we submit our last reports…

It’s amazing how quickly we get used to extraordinary things. Upon arriving at the station a month ago, we were facing our dream, facing what we had worked so hard to attain, and for that reason we felt invincible. Over the past weeks, we have dodged storms, lived under Mars’ atmosphere, played by his rules, encountered obstacles, and sometimes unfavorable odds. It is difficult now to look back and see what we have accomplished: it all still feels part of our day to day, of routine operations, and the tiny steps along the way don’t yet add up to the monumental leap we have taken. But in a little while, perhaps no later than tonight, we will feel even more invincible than on Sol 1: the scientific results we have produced and the friendships we have formed will appear, evidently, before our eyes.

At 5 pm, we opened both airlock doors, feeling the wind on our faces for the first time since we left Earth. If there had been water and canals in the Utah Desert, we might have seen some Martian traits showing through our very Earthly faces, smiling back at us.

EVA Report – March 10th

Crew 275 EVA 22 Report 18Feb2023
EVA # 22

Author: Quentin Royer (Crew Engineer)
Purpose of EVA: Retrieval of all the atmospheric instruments

Start time: 09:07
End time: 10:43

The 3 crewmembers drove with the rovers and arrived to destination at 9:17. They unloaded the rovers and headed towards the instruments. The EVA Crew started with the meteorological mast. No problem was encountered during its dismantlement, and it was safely packed.
The EVA Crew then began to dismantle the MegaAres antenna, and first retrieved the two hemispheres. As they were very light, and as there were high winds, the crew decided to bring the equipment which were already dismantled back to the Hab.
The rovers were loaded, and the crew departed the atmospheric instruments’ location at 9:50.
Then, the EVA Crew unloaded the rovers and put the equipment in the main airlock, for retrieval by the rest of the Crew located in the Hab.
The EVA Crew was back on site at 9:54. The disassembled the rest of the MegaAres antenna with no problem. Then, the LOAC was retrieved and packed.
The EVA Crew was back at the Hab at 10:34.

All the objectives of the EVA were completed.

Destination: Between the Hab and Marble Ritual, near the road

Coordinates (use UTM WGS 84): N4251000 E518500

EVA Participants :
• EVA leader: Quentin Royer (Crew Engineer)
• Alexandre Vinas (Crew Astronomer)
• Corentin Senaux (HSO)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual

Mode of travel: Driving, even if the destination was less than 1 km away, the crew needed to transport heavy instruments that also take up a lot of space.

GreenHab Report – March 10th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 69.5°F

Average humidity: 21%

Hours of supplemental light: 0 hour

Daily water usage for crops: 2.64 gallons (from the blue tank) & 10.57 gallons (from the aquaponics system)

Daily water usage for research and/or other purposes: 0gallons

Water in the blue tank: 44.23 gallons

Time(s) of watering for crops: 11:10

Changes to crops: N/A

Narrative: Today was the last day of the simulation, I then had to stop my experiments. I used the clean part of my aquaponics water to water plants and crops so as to waste as little water as possible.

Plants and crops look good; however, it seems important to keep an eye on the spinach as it seems to become a little bit soft. It is sad to witness that at the end of my mission…

However, I can’t tell you how delightful it is to see beans growing since I planted these, same for radishes, and cucumbers thriving.

Thus, overall, everything is fine, and future crews will be welcomed with greens and carrots as I saw!

It is with great pleasure that I worked in this environment for 4 weeks and a greater pleasure that I think about future crews working there.

Harvest: (From my aquaponics experiment) 146 g of iceberg salad, 125 g of ‘salad bowl’ (lettuce), 50 g of ‘salad blend’ (lettuce), 47 g of kale, 33 g of spinach, 18 g of basil

Support/supplies needed: As I said in the narrative, spinach seems to start to struggle, it would be nice to keep an eye on them.

As suggested by Donald, we will let you keep the fish, they are in a white bucket next to the tomato plants. Can you confirm that you will take them so that we know for sure that we do not need to take them with us?

Operations Report – March 10th

Crew 275 Operations Report 10Mar2023
SOL: 26

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• Robotic observatory

Notes on non-nominal systems: None.

Spirit rover used: No
• Hours: 210.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 114.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 222.0
• Beginning charge: 100%
• Ending charge: 90%
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 256.3
• Beginning charge: 100%
• Ending charge: 85%
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. An emergency depressurization protocol was simulated, and an emergency EVA was conducted, with Mission Support’s approval.
• WATER USE: 12.87
• Water: 317.24 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Starlink (MDRSCampus-Guest) is used during COMMS and during the day for emails and to handle the medical emergency. MDRScampus is used for experimental purpose.

Summary of suits and radios: One of the metal wires linking the collar to spacesuit n°4 broke. We replaced it with a small metal wire, as there doesn’t seem to be any long metal wires left.

Summary of GreenHab operations: The Crew botanist conducted routine operations and emptied the aquaponics system.
• WATER USE: 2.64 gallons
• Heater: On
• Supplemental light: No
• Harvest: 146 g of Iceberg salad – 125 g of “salad bowl” – 50 g of salad “salad blend” – 47 g of kale – 33 g of spinach – 18 g of basil

Summary of ScienceDome operations: No operations were conducted in the ScienceDome today.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: The robotic observatory is non-nominal.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support:
• The lamp located on the right when entering the ScienceDome is missing a light bulb.

Sol Summary Report – March 10th

Crew 275 Sol Summary Report 10Mar2023

Sol: 26

Summary Title: It’s not over until it’s over… but now it’s over!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

For our last EVA, we had planned four hours to disassemble all the atmospheric instruments located at a site close to the MDRS. We wanted to make sure that no adverse events could happen and that the EVA crew would have enough time for all the operations needed. But they were very efficient, and did not meet any problem. As a result, everything was done in less than two hours, even though the high winds did not make the operations as easy as they could have been.

The aquaponics system has been also dismantled by our GreenHab Officer. He took this opportunity to harvest the crops that grew there during our stay. The fish are still in their bucket, waiting for Donald to retrieve them. Please keep us updated, if he is not able to retrieve them, so that we are aware of it and can bring them back to the store where we purchased them.

Once we were all gathered in the Hab, we used our last hours there to complete all the protocols remaining for our datasets to be complete, which was successfully done. Another emergency drill was performed with the help of AI4U, our assistant based on an artificial intelligence. When the alarm rang at 13:55, our Crew Journalist again took the lead of the operations. One team was sent to go through all the buildings and detect the one causing the non-nominal status. As they were in the Science Dome, they simulated another issue, as a crewmember pretended to faint. While the HSO took care of this crewmember, the others prepared for an emergency EVA. They quickly, but safely, left the Hab and fixed a fictive leak located in the GreenHab. The drill took 30 minutes, which is exactly the time limit that was fixed.

On their way back from the emergency EVA, the EVA crewmembers found some items from our burnable garbage flying around. They managed to bring some of them back to the Hab. We saw more of them along the pressurized corridors and tried to call Mission Support from 15:00. When we managed to reach them, we also took this opportunity to inform them about our wish to exit the sim, which was done a few minutes later, at 17:00.

Look Ahead Plan:

Now that we are out of simulation, we will complete the checkout list as well as the food inventory. We will also start gathering all our personal belongings and scientific material. At what time can the CrewCar be placed in front of the Hab so that we can start loading it? We plan to leave tomorrow at 12:00. We will fill the tank of the CrewCar in Grand Junction, park it at the Day’s Inn, and give the keys at the front desk.

Some water was spilled on the computer of the Commander. Now it does not turn on anymore. This computer contained the Final Mission Science Report as well as the Final Mission Summary. Would it be possible to send these to you tomorrow instead, so that we have time to write them again?

Anomalies in work: None.

Weather: Windy, low 3°C/37°F, high 12°C/54°F.

Crew Physical Status: One crewmember recovering from surgery after a broken collarbone.

EVA: EVA #22 dedicated to the disassembling of the atmospheric instruments conducted successfully. EVA #23 (emergency EVA) dedicated to a drill of an emergency protocol in the vicinity of the Hab also conducted successfully.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site